What Is The Absolute Difference In [H+] Between 2 Aqueous Solutions, One Of PH 2.0 And One Of PH 3.0? What is the [OH-] concentration of the solution of pH 2.0?


1
Expert's answer
2021-03-30T04:57:40-0400

As we know that,


"\boxed{

P^H=-log[H^+]}"

Therefore,

as given PH=2 and PH=3


"2=-log[H^+]\\10^{-2}=[H^+]" ...ans


and


"3=-log[H^+]\\10^{-3}=[H^+]" .......ans


So absolute difference between 1 and 2 is

"\boxed{10^{-2}-10^{-3}=0.009}"



Further for concentration of [OH-] of PH= 2

We know that,

"\boxed{[OH-][H+]=10^{-14}}"


so from this ,


[OH-][10-2]=10-14

[OH-]=10-12 ,....answer